Not sure on this word problem
Adult tickets for a play cost $20 and child tickets cost $16. If there were 27 people at a performance and the theater collected $464 from ticket sales, how many adults and how many children attended the play?

Answer by scott8148(5880)

a+c=27 ... a=27-c ... a(20)+c(16)=464 ... (27-c)20+c(16)=464 ... 540-20c+16c=464 ... -4c=-76 ... c=19

a=27-(19) ... a=8